The equation for Perimeter is P = 2L+2W. Solve for W. *

From your answer in question 7, find the width of a rectangle that has a perimeter of 10 and a length of 3.

Answer by algebriac(10)

The equation for Perimeter is P = 2L+2W.
Hence 2W= P- 2L
W= (P-2L)/2
Now given P=10 and L=3
W= (10-(2*3))/2
=(10-6)/2
= 4/2
W= 2